OpenStudy (anonymous):

Solve this system. Please state answer as an ordered pair. 2x-y=-1 3x+5y=18

OpenStudy (anonymous):

You have to use substitution or elimination

OpenStudy (anonymous):

I'll just use elimination

OpenStudy (anonymous):

so, Would it look like 2( 6 - 1.6y ) -y = -1 I don't think i got it right..

OpenStudy (anonymous):

No that's not it.

OpenStudy (anonymous):

I firgured..

OpenStudy (anonymous):

*figured

OpenStudy (anonymous):

Never mind i'll use substitution.

OpenStudy (anonymous):

2x-y=-1 -2x-2x -y=-2x-1(-1) y=2x+1 Now substitute y=2x+1 into 3x+5y=18. 3x+5(2x+1)=18 3x+10x+5=18 13x=13 x=1

OpenStudy (anonymous):

Now substitute x=1 into 2x-y=-1 2(1)-y=-1 2-y=-1 -y=-3 y=3

OpenStudy (anonymous):

The ordered pair is (1,3)

OpenStudy (anonymous):

does that help @Keridavis

OpenStudy (anonymous):

Yes, thank you